Find the equation for the line that passes through (2,-2) and that is PERPENDICULAR to the line with the equation (y-3)=-1/4(x+3)
Answer by lwsshak3(11628)

Find the equation for the line that passes through (2,-2) and that is PERPENDICULAR to the line with the equation (y-3)=-1/4(x+3)
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
(y-3)=-1/4(x+3)
y-3=-(1/4)x-3/4
y=-(1/4)x-3/4+3
y=-(1/4)x-15/4
slope=-1/4
..
Slope of line perpendicular to given line=4 (negative reciprocal)
equation: y=4x+b
solving for b using given coordinates(2,-2)
-2=4*2+b
b=-10
Equation of perpendicular line:
y=4x-10